### v1

# Spatially-Embedded Lens Visualization: A Design Space

Abstract
Lens visualization has been a prominent research area in the visualization community, fueled by the continuous need to mitigate visual clutter and occlusion resulting from the increase in data volume. Interactive lenses for spatial data, particularly, challenge designers to conceive design strategies to support the analysis of high-density, multifaceted data with spatial referents. Despite their relevance, there is a lack of systematic understanding regarding the various design elements that compose spatially-embedded lens visualizations. To fill in this gap, we unify these components under a common hood in the form of a design space, which we propose in this paper. Building our knowledge on top of the initial insights gained from Tominski et al.’s survey [57], we construct a design space spanning 7 dimensions through our analysis of 45 papers published in the visualization community over the past 15 years. We describe each design dimension through representative examples and examine the range of design choices available within each, discussing their benefits and pitfalls that affect lens performance and usability. In doing so, we offer a cohesive catalog of considerations for designers—both when examining existing lenses and when conceptualizing novel spatially-embedded lens visualizations. We conclude by shedding light on regions of the design space that remain largely understudied, revealing open opportunities for future research.
keywords:
Focus+Context, Lenses, Visualization, Design space, Taxonomy, SurveyK.6.1Management of Computing and Information SystemsProject and People ManagementLife Cycle;
[…]
A collage of lens visualizations categorized across seven design dimensions.
Position, orientation & scale: interactive (, ) and semi-automated ().
Shape: fixed (, ), dynamic user-driven (, ) and data-driven ().
Dimensionality: 2D (, ), 2.5D (, ), and 3D (, ).
Effect scope: separate view (, ), lens interior (, ) and exterior (, ).
Effect imagery: 2D (, ), 3D (, ), and Decal ().
Effect encoding: thematic (, ) and spatially-based (, ).
Viewpoint Dependency: independent (, ) and view-dependent (, ).
[…]
Introduction
As data size is constantly increasing, visualization techniques have to deal with the problem of visual clutter and occlusion—occlusion hides important information from the viewer, whereas clutter shows too much visual information and makes the visualization more complex for the viewer to understand and gain insight from it. The high data density and dimensionality associated with spatial visualizations, particularly, pose significant challenges in encoding all relevant information into a single visual imagery. To overcome the challenges of clutter and occlusion, the visualization community has embraced the idea of interactive lenses—a family of focus+context techniques that provide on demand an alternative visual representation of a focus area of the screen, while the remaining regions convey context [57].
In the last decades, numerous lenses have been proposed in the visualization research, with a growing number in the field of immersive analytics which leverages immersive technologies for visual data analysis [57, 15]. Despite their relevance, there is still few work that provides a more systematic knowledge on the question of what are design components and choices to be considered when constructing visualization lenses? This is particularly true in the case of lenses that address the challenges encountered in spatial visualizations.
Contribution To fill in this gap, we contribute in this paper a design space of interactive lenses for spatial data, namely spatially-embedded lens visualizations. Through our survey and analysis of 45 exemplary lenses published over the past 15 years in the visualization literature, we systematically formulate 7 design dimensions that capture the salient aspects of interactive lenses operating within spatial environments—contemplating both conventional and more immersive spaces.
Outline In Section 1, we introduce the core concepts of lenses. In Section 2, we proceed to describe our design space by explaining each design dimension with examples, and outlining the available design choices within each dimension along with their trade-offs that influence lens performance or usability. In Section 3, we demonstrate the applicability and generative capacity of our conceptual schema through a detailed design use case. Lastly, in Section 4, we leverage our design space to identify and discuss as-of-yet understudied design areas that reflect opportunities for future research.
1 Lens Conceptual Model & Definition
Conceptual model Tominski et al. define a conceptual model of visualization lenses built upon the well-established visualization pipeline [57], as illustrated in Fig. 1. Taking this as a basis, the lens pipeline has three main stages. First, the lens selection () captures what part of the data is to be affected by the lens. Second, the lens function () is the operator that defines how the lens modifies the visualization, and may depend on parameters that control the lens effect—e.g. a zoom factor of a magnification lens, or an alpha or threshold value of a filtering lens. Lastly, the join () stage joins the lens effect with the base visualization. We incorporate this conceptual model along with its terminology throughout our manuscript.
The definition of visualization lenses originates from the idea of a physical lens moved above an area of interest to temporarily magnify the visibility of details. However, this traditional interpretation has been significantly broadened in the context of visualization [57]. Beyond solely magnifying visual elements, lenses have been extended to include a range of functions for manipulating the visualization. In general, a lens function can modify a base visualization by altering existing content, suppressing irrelevant content, augmenting with new content, or even a combination thereof—see Fig. 2.
Furthermore, although lens effects are often transient by temporarily manipulating the view of focus areas, they may also be used to permanently modify the visualization—e.g. copy a visual element located under the lens boundary and paste copies of it in other parts of the visualization [8, 52]. Finally, lens functions are interactively parametrizable. This means that users may interact with the lens to adjust parameters that control the lens effect while carrying out visual data exploration and analysis—e.g. the zoom factor or even the mathematical function specifying the magnification distortion.
Definition Considering the previously discussed aspects, we consider the following definition of visualization lenses: an interactively parametrizable spatial selection according to which a base visualization is modified to provide an alternative visual representation of the data in focus 111Similar to the definition of Tominski et al. [57].
2 Design Space
Although numerous visualization lenses have been introduced over the past decades, the question of what are design components and choices to be considered when constructing visualization lenses? remains understudied, as few work have attempted to establish a structured space that documents lens techniques by describing them in well-defined design elements. Establishing a design space is valuable since it allows us not only to systematically examine existing lenses but also to reason about new lens-based techniques. This section attempts to define such a design space of spatially-embedded visualization lenses.
An early taxonomy on lenses was proposed by Bier et al. [7], and later extended by Tominsky et al. [57]. The latter taxonomy encompasses both practical aspects that relate to lens usage, such as data types, user tasks, and display mediums, as well as conceptual aspects primarily concerning the visual design of lenses. Here, we aim to extend the existing conceptual schema as listed in Table 1. To this end, we discuss the Position, Orientation & Scale and Shape dimensions, which have been previously defined as properties driving lens selection [57]. The Effect Scope dimension is defined as the area within the visualization where the lens effect is applied, and is a refinement of the Effect Extent class from the previous conceptual schema. The remaining dimensions—namely, Dimensionality, Effect Imagery, Effect Encoding, and Viewpoint Dependency, have not been formally defined in the literature so far. We explain each of the aforementioned design dimensions with examples, and outline the set of possible design choices within each dimension along with their trade-offs that influence lens performance or usability.
Effect Class: | Suppress, alter, enrich |
Effect Extent: | Lens interior, side effects, separate view |
Adjustability: | None, interactive, self-adjusting |
Selection Stage (): | DS, DT , VA, V |
Join Stage (): | DS, DT , VA, V |
Pos, Ori & Scale: | Interactive, semi-automated |
Shape: | Fixed, user- or data-driven dynamic |
Dimensionality: | 2D, 2.5D, 3D |
Effect Scope: | Interior, exterior, separate view |
Effect Imagery: | 2D, 3D, Decal |
Effect Encoding: | Thematic, spatially-based |
Viewpoint Dependency: | Invariant, view-dependent |
Sampling criteria To collect relevant papers for our review we established three inclusion criteria:
. Papers published in the past years, from to .
. Papers presenting techniques that fit into our definition of visualization lens—see Section 1, which included techniques that call themselves lenses or have lens-like characteristics.
. Papers demonstrating lens techniques that address spatial data in either 2D or 3D spaces—by spatial data we mean data with inherent width, height, or depth, where the relative positioning, length, shape, etc. are meaningful for users. Although our discussion in this paper focuses on 3D data, our decision to consider both spatial 2D and 3D visualizations was based on their shared need to account for the spatial nature of the data. Consequently, they share concerns and constraints when designing lenses—e.g. accurate spatial perception, spatially coherent transitions between focus and context areas when applying lens effects, as well as occlusion and clutter management when representing multidimensional data with specific spatial referents. In doing so, we expanded our paper sample to facilitate extracting common lens design strategies and formulate our design space.
. Papers proposing lens-based visualizations in either conventional or more immersive environments. Our decision to include immersive lenses was aimed at forming a more comprehensive taxonomy that incorporates design aspects accounting for emerging displays, thereby contributing to the growing field of immersive analytics.
Methodology To find relevant papers for our review we used a snowball sampling technique starting with the seminal paper by Bier et al. [8], which first introduced the concept of interactive lenses for visualization. From this starting point, we recursively scanned references for further papers about lenses, and we stopped searching and collecting papers once this process reached saturation and no new relevant papers were identified. From the collected papers, we derived our design space through open coding, based on consensus among all authors. The resulting design codes and coded papers can be seen in Fig. 3.
Drawing from our literature sampling and coding, we analyzed papers and derived a design space with design dimensions, systematically describing many design components and choices to be considered when constructing visualization lenses for spatial data: Position, Orientation & Scale, Shape, Dimensionality, Effect Scope, Effect Encoding, Effect Imagery, and Viewpoint Dependency. Although not exhaustive, our design space forms a systematic understanding of available design components, choices, and their trade-offs—summarized in Table 2—in order to guide the development of new lens visualizations. The following subsections catalog our design space by explaining each design dimension along with relevant examples.
2.1 Position, Orientation & Scale
Because of the tight interrelation between position, orientation, and scale, we consider them together within a single design dimension. These lens parameters primarily specify the lens selection.
Most lens techniques supply users full interactivity for setting the lens position, orientation, and scale.
Tominski et. al proposed a circular lens where users interactively set its position and radius, delimiting the lens selection over spatiotemporal data on 2D maps [58].
Lens orientation, along with position and scale, can influence focus selection.
Dumas et al. proposed an angular-based lens technique for selecting curves within 2D visualizations—e.g. spatiotemporal movement data on maps. The lens’ circular selection is specified not only by its position and diameter but also by direction and angular tolerance [19].
Orientation is often overlooked in 2D spaces but becomes more essential when working with 3D data.
By adjusting orientation, the lens can align its geometry with the underlying 3D visualization—e.g. aligning a spatially-tracked mirroring lens with intertwined blood vessels [6].
Although most surveyed lenses are fully interactive, some rely on (semi-)automated control based on the underlying data [18, 1, 3].
Semi-automation may be useful particularly in scenarios where domain knowledge is encoded into the automation logic, guiding users towards relevant data features—e.g. flow patterns.
It can also enhance precision in spatially-constrained visualizations or in tasks requiring more fine-grained lens manipulations, like route-following.
To illustrate, Alvina et al. introduced a magnification lens to reveal details while users steer along map routes [1].
Although it is ideal to keep the lens center close to the route—since fisheye distortions aggravates with distance from the center, traditional magnification lenses often fail and let paths “slip off” the lens.
To address this, the lens automatically adjusts its position its position based on the route geometry, making routes easier to follow.
Although semi-automated lenses can reduce user effort and enhance precision, they may reduce user agency if the underlying algorithms misalign with users’ intent. Additionally, they may highlight irrelevant features if heuristics fail to generalize across datasets.
2.2 Shape
The shape of a lens refers to the geometry used for the lens selection, and can be classified as either fixed or dynamic—see Fig. 4.
Most lenses proposed in the literature are of fixed shape.
While they typically use simple, primitive forms—e.g. circular, rectangular, box [42, 28, 23], there are no restrictions preventing rigid, non-standard shapes. For instance, Trapp et al. proposed a lens to display different levels of structural abstraction in 3D city models [61]. The lens technique uses predefined 3D models as custom lens shapes, created using external modeling software.
Fixed lens shapes typically offer simplicity and familiarity. They are easy to implement and interact with, which can lower the learning curve. However, they may not adequately conform to irregular data structures, resulting in poor coverage of relevant features, inclusion of irrelevant ones, and mental burden as users themselves must disambiguate them.
Some lens techniques adapt their shape dynamically, either by the user or by the underlying data
In the former case, lens shapes are typically set by users through sketching, brushing, or lassoing (Fig. 4–right) [48, 11, 24].
Brosz et al. proposed a lens-like technique where shapes can be sketched over 2D maps [11]. Users draw an origin shape that matches a target feature, such as a route, and a destination shape. The visual content within the origin is then morphed to fit the destination form.
Kister et al. proposed a body-centric lens whose shape matches the body silhouette, like a shadow cast by the body onto the screen [32].
User-driven dynamic shapes provide greater versatility and precision for selecting irregular features, better reflecting user intent. The trade-off is increased interaction effort—e.g. manual drawing in dense 3D spaces can become cumbersome.
For the latter case, data-driven dynamic lenses automatically adapt their shape to match data features—e.g. boundaries, clusters, paths.
Pindat et al. proposed a 2D magnification lens that implicitly adapts to the geometry of a focus object underlying the mouse cursor—e.g. a topographic path or area.
This shape morphing aims to provide less-distorted magnified images compared to traditional radial fisheye lenses [45].
Ion et al. introduced magnification lenses with dynamic shapes to maintain focus of moving data targets on a map that become off-view of the lens boundary (Fig. 4–middle) [25].
Data-driven automation can reduce user effort, enable automated feature discovery, and ensure lenses align meaningfully with the data—thus improving interpretability. However, these lenses may behave unpredictably when the data is noisy or lacks clear structure, making it difficult for users to understand or control the logic behind the autogenerated shape configuration. Therefore, implementing such behavior requires careful data preprocessing and logic calibration or transparency to ensure reliability.
2.3 Dimensionality
The dimensionality is classified either as 2D, 2.5D, or 3D—see Fig. 5, and relates to the other geometric parameters of lenses that drive the lens selection—i.e. position, orientation, scale, and shape.
2D lenses are the most commonly used and have polygonal shapes—-e.g. circular or square, which are placed and manipulated in screen space.
For instance, Tim and Elmqvist proposed the use of physical lenses made of thin, transparent paper sheets on top of tabletop surfaces [31]. Their physicality affords them to be intuitively overlapped, creating composite lens effects on the underlying visualization.
Although 2D lenses have been used in 3D visualizations [40, 32, 24, 13], their 2D nature constrains the lens manipulations to the view plane and leads to two primary limitations when handling 3D data: an inability to carry out a 3D selection, and an absence of spatial correlation between the 2D alignment of the lens and the underlying 3D visualization.
2.5D lenses consist of 2D lenses embedded in 3D spaces, thereby augmenting lens manipulations to 3D [42, 33, 23, 6]. Although this alleviates the issue of spatial correlation, it increases interaction effort to place and align the lenses according to the underlying data. In addition, the flat nature of 2.5D lenses constraint the spatial selection to only a slice of the 3D data. To illustrate, Spindler et al. introduced a spatially aware, handheld paper sheet lens to improve lens manipulations in 3D spaces above tabletop displays [53]—e.g. users could arbitrarily slice an MRT scan of a human torso in different poses.
3D lenses consist of 3D volumes—e.g. a sphere,
which enable volumetric spatial selections of the 3D data [28, 61, 9]. Due to their three-dimensional nature, the lens manipulations are prone to suffer from interaction effort similar to 2.5D lenses. Additionally, such lenses may not closely align with surface geometries due to their typically fixed shapes; this can lead to perceptual issues such as the lens itself occluding parts of visualization—especially those of intricate geometry. Viega et al. first introduced 3D lenses to limit the lens effect to a finite subvolume of interest [63]. The lens effect is built from the intersection of GPU clipping planes for revealing hidden structures of 3D datasets.
2.4 Effect Scope
The lens effect scope refers to the area within the visualization where the lens applies its effect, which can be classified as lens interior, lens exterior, or separate view—see Fig. 6.
Traditionally, a lens function only applies to visual elements confined inside its shape bounds—-that is, the lens effect scope is the lens
interior [42, 24, 3].
Restricting the lens effect to its interior maintains a clear distinction between focus and context. It also localizes visual changes, minimizing disruption in the surrounding visualization. To illustrate, Hurter et al. presented a lens that selects a specific spatial and attribute-related data range. The lens preserves the data in focus but continuously deforms in-lens data outside the selection range, moving it toward the lens border (Fig. 6–right) [24].
Although interior effects offer localized focus+context visualization, it may obscure broader patterns or connections extending beyond the lens boundary—especially in tasks requiring spatial continuitiy, like tracing trajectories.
A few lens effects impact visual elements located not only inside the lens boundary but also beyond it; therefore, the effect scope is considered to be the lens exterior [41, 50, 19]. In such cases, the lens boundary is primarily used to define a selection based on a local behavior of interest in the underlying data—e.g. to select trajectories based on a local directionality.
To illustrate, Mota et al. introduced a lens that highlights streamlines passing through the lens boundary and having slopes locally oriented approximately in the same direction as the lens (Fig. 5–right) [41].
Tong et al. proposed deformation-based lenses to clear the line of sight between the viewer and 3D glyphs in focus [60]. The lenses remove occluding glyphs situated within the lens boundary by displacing them aside in the surrounding context.
Exterior effects are also useful when users wish to explore patterns emerging across multiple foci—by composing two or more lenses to build cascaded selections or effects across spatial regions [35].
For instance, Scheepens et al. proposed a lens for trajectories in 2D visualizations (Fig. 6–middle) [50]. The technique supports composing multiple lenses to enable compound selections of more convoluted traffic flows of interest.
Exterior-scope lenses, however, can weaken spatial associations between the lens and its effect, creating ambiguity in attributing effects to their originating lenses.
Some lenses render their effects in independent visual spaces by generating entirely new views separate from their source selection.
They commonly showcase either a duplicate of the visual content enriched with additional elements, or an alternative visual encoding that may not be suitable in the originating view.
This effect scope is referred to as separate view [46, 10, 50, 27, 58].
Lenses whose effects are displayed across separate views enable side-by-side comparisons; however, careful layout design is needed as to avoid becoming cognitive demanding by requiring users to mentally relate spatially-detached views.
Therefore, a key design consideration is how to spatially relate the separate views to their corresponding lenses—e.g. via visual links or color coding [50].
Karnick et al. proposed a lens that uses a layout algorithm to determine optimal placement of focused views near referent locations along a selected geographical path in a 2D map [29]. The design priorities of the layout algorithm were to prevent the lens views from occluding the base map visualization and to maintain proximity to their associated points of interest.
Furthermore, having separate views enables the definition of cascading effects, where each lens presents a distinct data transformation while preserving the history of prior lens stages.
Javed et al. proposed a lens-like technique that progressively constructs a hierarchical tree view from the user-defined, cascaded focus regions in a 2D map (Fig. 6–left) [27].
Subsequent tree levels display views with higher magnification degrees, and the layout algorithm concentrates on forming clear parent-child relations to aid navigation along the cascaded hierarchy and on positioning views at a given tree level side by side to enable easier comparisons.
2.5 Effect Imagery
The effect imagery refers to the visual image generated by the lens function, and can be classified as 2D , 3D, or Decal images—see Fig. 7.
In 2D lens images the imagery generated as the lens effect consists of projected images, which are rendered either directly onto the view plane [23, 24] or onto a separate render texture to display the lens image as a separate view [42, 46, 6, 5].
As an example, Pindat et al. proposed a lens that renders its projected image in a separate viewport, which users can move and rescale. The 2D lens image contains the same visual elements rendered within the lens boundary, but with a different magnification level and viewpoint [46].
A limitation of 2D images is providing only a monocular view, which may result in reduced spatial perception of depth and shape of the 3D data visualization. Therefore, an important design consideration is whether and how to use such images for visualization tasks that rely on shape and depth information—e.g. identifying isolated blood vessel pathways may be challenging, as they could appear ambiguous and not clearly discernible from a static projected image due to overlaps, occlusion, and the lack of depth perception in a 2D lens image [6]. In this sense, a view-dependent 2D image may assist restoring the lost spatial perception, as motion parallax effects manifest when projections are continuously rendered in real time [42, 6].
3D images consist of volumetric visual images, possibly rendered as a separate volume and enriched with additional visual elements [9, 28, 5]—similar to a world-in-miniature [54].
As an example, Johnson et al. presented a technique for exploring 4D cardiac data using a global-to-local approach. Each lens selection generates a magnified 3D duplicate displayed in a separate view, and these views are automatically arranged in a juxtaposed grid. This enables visual comparison of key features across scales and points of view while minimizing irrelevant regions.
Due to its volumetric nature, 3D images preserve spatial perception better compared to 2D projected images. However, rendering and interacting with 3D imagery can be significantly more computationally costly than with 2D imagery.
This is particularly concerning as immersive visualizations demand both high resolution and high frame rates, posing a greater challenge compared to visualizations on conventional displays, where lower frames rates and intermittent pauses for computations or data loading are more tolerable.
To mitigate latency challenges, Johnson et al. proposed a combination of data sampling strategies and clipped volume rendering to display volumetric images at interactive frame rates specifically for use in virtual reality [28]. Borst et al. proposed a data tiling technique to speed up rendering of multiple volumetric images of focus regions on topographic surfaces [9].
Decal images conform to surface areas of interest. This conforming behavior not only establishes a strong spatial correlation between the lens image and its referent location but also mitigates perceptual issues, such as the lens image itself occluding parts of the base visualization—particularly those of intricate geometry.
To illustrate, Rocha et al. proposed decal lenses for multivariate data visualization on surfaces [48]. The lens image is built from the intersection of a sphere with a surface. The sphere itself does not define the area where the lens effect appears; rather, this area corresponds to the portion of the surface that lies within the sphere. This leads to a lens image that conforms to the surface geometry, similar to a decal.
Mota et al. later proposed a hybrid lens that acts on different data types that commonly co-exist in 3D visualizations: streamlines and surfaces [41]. The lens integrates two categories of lense images—3D and Decal—to become a versatile lens that applies different effects according to the underlying data type in focus—non-surface and surface data, respectively.
Dimensionality | Effect Encoding | Effect Imagery | |||||
2D | 2.5D | 3D | Thematic | Spatial | 2D Image | ||
Standard 2D interaction* | Spatial interaction* | Volumetric selection* | Higher-level reasoning* | Spatial reasoning* | Familiar representation* | ||
Simpler lens layering* | Slice-only selection | Higher interaction effort | Abstract-physical gap | Anchored visual cues* | Lower rendering cost* | ||
Lacks spatial correlation | Misaligned, rigid selection | Relies on user decoding | Low spatial perception | ||||
Prone to clutter |
Shape | Viewpoint Dependency | ||||||
Fixed | DYN (User-driven) | DYN (Data-driven) | Dependent | Invariant | 3D Image | ||
Easy to use* | Custom shape control* | Aids feature discovery* | Responsive to user intent* | Stable across view changes* | Higher depth perception* | ||
Easy to implement* | Good shape–feature fit* | Tigher shape-feature fit* | Intuitive for immersion | Ignores view awareness | Latency risk | ||
Poor feature coverage | Manual interaction effort | Reduced user effort* | Unintended effect changes | ||||
Data-sensitive fragility | |||||||
Opaque logic |
Effect Scope | Position, Orientation & Scale | ||||||
Interior | Exterior | Separate View | Interactive | Semi-automated | Decal Image | ||
Focus-context separation* | Reveals peripheral patterns* | Supports cascading effects* | Full control* | Data-driven targeting* | Strong spatial anchoring* | ||
Localized effect* | Enables multi-foci effects* | Persistent lens history* | Fosters exploration* | Aids precision* | Less lens-induced clutter* | ||
Hides broader links | Weaker lens-effect binding | Cross-view burden | Demands spatial dexterity | Misleads if heuristics fail | Fragile on uneven surfaces | ||
Generalization liability |
2.6 Effect Encoding
The lens effect enconding can be classified either as thematic or spatially-based, as shown in Fig. 8.
Thematic encoding draws from information visualization and refers to abstract representations of the lens’ data selection—e.g. linear and radial histograms (Fig. 8–left) [58, 40, 36].
Hurter et al. proposed a dual-layout lens that, within its boundary, rearranges pixels representing a 3D model’s color-coded cells by transforming them from a Cartesian layout into a polar histogram layout [24].
Scheepens et al. proposed a lens for map-based visualizations that generates thematic images as separate views to show traffic flow over time [50]. Users can reposition and stack the views, triggering Boolean operations—e.g. computing the difference between the originating views.
Conversely, spatially-based encoding
draws from scientific visualization and concerns physically-based data representations—e.g. glyphs, contours, and isosurfaces (Fig. 8–right) [48, 60, 23].
Gasteiger et al. presented a lens to support cerebral aneurysm analysis, where analysts must correlate surface attributes with and adjacent blood flow—e.g. the wall shear stress and underlying inflow jet [23]. This technique, places seed points within the lens boundary, generating illustrative color-coded streamlines, while surface attributes are depicted as isosurfaces or saturation-coded contour lines.
Rocha et al. presented a decal lens to allows users to locally switch attributes on arbitrary surfaces, obtaining alternative data representations—e.g. glyphs [48].
While both thematic and spatial mappings aim to represent data, they are shaped by different driving forces and usage purposes [22]. Thematic mapping emphasizes higher-level reasoning. It has been used to analyze multiple attributes simultaneously, and to uncover patterns not directly visible in spatial data—e.g. correlations, trends, outliers, or temporal cycles [44, 55, 2]. However, abstraction may introduce cognitive load by obscuring the underlying spatial structure of the data, requiring users to mentally relate abstract visuals back to the physical space. In contrast, spatial mapping emphasizes spatial reasoning by preserving the data’s geometric and topological structure. This increases users’ spatial awareness, facilitating the correlation of attribute data with physical referents. However, spatial encodings—e.g. contours, glyphs, or color mapping, heavily rely on users to interpret visual representations. With less abstraction to guide and summarize meaning, users themselves must perceive shapes, spatial relations, patterns, and other visual cues. They are also prone to clutter, which reduces readability—particularly in dense or 3D environments. Thus, careful design is required to balance spatial fidelity with visual clarity.
2.7 Viewpoint Dependency
Lenses can be classified either as dependent or invariant from users’ viewpoint, as shown in Fig. 9. This dependence can involve either the lens selection, function, effect, or even a combination thereof.
The degree to which a visualization depends on the user’s viewpoint is a key design consideration—especially in immersive spaces,where users can move physically [47, 26]. This is particularly relevant for immersive focus+context visualization, where the focus reflects user’s current interest. However, physical locomotion can raise visibility and perceptual issues in the focused visualization: visual encodings may become imperceptible with increasing distance (leading to change blindness), distort under acute viewing angles, or even obstruct other foci distributed in the space.
Although most papers do not consider viewpoint in their designs, view-sensitive lenses afford fluid and intuitive interactions by dynamically adapting to users’ physical navigation or gestures—e.g. automatically expanding when users move away, and revealing fine-grained information when in close proximity (Fig. 9–right) [32]. These dynamic responses allow lenses to better anticipate and align with user intent, likely reducing the need for manual input. However, design considerations are needed to avoid unstable and disorienting effects caused by unintentional movements—e.g. to use smoothing or “lock-in” mechanisms.
For instance, a handheld mirroring lens should account for minor involuntary hand movements, common in immersive environments, to prevent jitter in the reflective effect [5, 6, 56].
Nearly all lenses are viewpoint independent, ensuring stable lens images across view changes. However, this comes at the cost of ignoring view awareness—e.g. lenses may appear occluded, misaligned, or illegible from certain distances or angles. As a result, users must compensate through manual input, increasing interaction effort and possibly disrupting the analysis workflow—see Fig. 9–left. Ebert et al. introduced tangible ring objects that function as lenses placed on tabletop displays, where users physically manipulate them to update their associated positions [20]. Käser et al. proposed magnifying lenses, where users control both position and scale with their fingers [30].
3 Generative Use Case
This section demonstrates the descriptive and generative utility of our design space by illustrating a synthetic case analysis that characterizes three spatially-embedded visualization lenses as functions of our dimensions. This use case was motivated by the observed scarcity of research on composite lenses. Composite lenses recursively combines multiple lens effects, enabling users to activate and deactivate individual effects as necessary. Additionally, our ideations for this case stems from our long-term collaboration with domain experts as well as prior experience in developing visualizations applied to reservoir engineering.
3.1 Design Case: Composite Lenses
In the oil and gas domain, a key visual analysis task is correlating variations in the flow behavior with attribute changes occurring in intertwined geological channels. To illustrate, one may imagine a scenario where water is injected from an injecting well to displace and produce oil in the producing well , as depicted in Fig. 10–a.
To address the inherent occlusion caused by the volumetric nature of reservoir grids, an engineer activates an interactive cutaway lens. This 3D lens has the form of a frustum centered at the user’s point of view—see Fig. 10–b, and continuously clips all reservoir grid cells located within its interior. This clipping behavior produces a projected cutaway view of the reservoir that dynamically changes according to the user’s head and whole-body movements, thereby freeing their hands to perform other analysis tasks—e.g. manipulating other lenses. In this way, the view-dependent cutaway lens enables the engineer to physically navigate and inspect the internal parts of the reservoir grid—i.e. the geo channels and water/oil flow paths.
Figure 10–c shows a cut region with two color-coded geo channels penetrated by the producer well —yellow and orange channels, and the underlying oil flow represented by red arrow glyphs. Near the well , the engineer observes a steady oil flow coming from the leftmost yellow channel, while the rightmost orange channel produces no oil. The engineer then proceeds along the orange channel towards the injecting well , examining this unexpected absence of flow. Eventually, the engineer reaches a location where the pathways intersect once again, with flow behavior akin to that near (Fig. 10–d).
To examine the local flow, the engineer activates an interactive, view-invariant multiview lens (Fig. 10–e). This circular 2.5D lens captures the current viewpoint within its area. It generates one or more color-coded, individual views—projected images of the same region showing different attributes that influence flow, like water saturation , permeability , and oil saturation (Fig. 10–f, g, h). By inspecting the multi-attribute lens images, the engineer observes high permeability and oil saturation in the orange channel, but also a thief zone with elevated water saturation . The engineer then suspects that the higher viscosity of oil compared to water causes the injected water from to preferentially flow through the thief zone, displacing only the lower-viscosity water and leaving oil behind.
To confirm this hypothesis, the engineer activates an interactive, view-agnostic thematic lens to inspect flow through two regions of interest within each geo channel. A spherical 3D lens selects the reservoir grid cells contained in each region, generating a thematic view shown inside a decal lens image. In regions with intricate geometry—e.g. sharp discontinuities or pronounced curvatures, a redesign might use proxy surfaces as simplified, smoother approximations for decal placement [48]. This decal-based design reduces clutter and strengthens the spatial correlation between the lens imagery and the reservoir region in focus (Fig. 10–i). The thematic view portrays a radial histogram of two temporal attributes for the selected grid cells: water (blue) and oil (red) flow rate variations (-axis) over time (-axis).
By inspecting the two lens images, the engineer observes that the leftmost thematic view depicts the expected flow behavior—i.e. similar rates of oil and water flow, as the injected water effectively sweeps oil from the yellow channel (Fig. 10–j). Conversely, the rightmost thematic image indicates increasing water flow rates alongside negligible oil flow, confirming the hypothesis that these anomalous flow behaviors result from the formation of the prior thief zone for the injected water throughout the orange channel (Fig. 10–k).
4 Discussion, Research Directions & Limitations
This section discusses open research directions for lens visualizations by examining underexplored areas in the design space—see Fig. 3.
Position, Orientation & Scale Our survey indicated that most lens techniques adopt full interactivity ( out of papers), thus relying entirely on users to manipulate the lens parameters. Beyond that, nearly all semi-automated techniques consider conventional 2D spaces rather than 3D immersive environments ( out of ). Therefore, further research on assisted POS could be particularly beneficial in immersive mediums, not only for guiding lens placement toward interesting data features but also for assisting in precise lens manipulation within 3D spaces—since prior work provides evidence that spatial interactions often suffer from precision issues due to the absence of physical surfaces that provide force feedback [4].
Furthermore, semi-automation could benefit immersive lenses with cascaded images displayed across various separate views by determining optimal layouts that assist logical linkage and sensemaking of information through spatial arrangements. As limited research has examined layout possibilities, open opportunities remain to explore novel semi-automated 3D spatial layouts and their trade-offs in immersive spaces, as they differ significantly from flat grid layouts used in conventional displays. To illustrate, a recent study found that users prefer 3D spherical cap layouts when assessing hierarchical lens-like images of maps with varied magnification levels in immersive spaces [49]. However, the study also revealed that users frequently rearrange the surrounding images during analysis tasks.
Shape The shape is a key factor in determining proper lens selection and its associated effect. As most lenses proposed in the past years are of fixed shape ( out of surveyed papers), users may have adapted to them and, as a result, must mentally compensate for their limitations. For instance, magnification lenses of static shapes may fail to provide clear and unambiguous magnifications of irregular regions of interest—either the lens is too small, requiring continuous viewpoint adjustment as the region does not fully fit in the focus, or it is too big, encompassing the entire focus region but at the expense of also including surrounding context. In the latter case, users must mentally segment the presented magnified imagery; however, mental burden may become overwhelming when the focus data features are sufficiently intricate, and when motion parallax is missing as seen in static 2D projection lens images.
As our survey revealed that lens shapes dynamically generated either by the user or by the underlying data is a widely underexplored area ( papers), future research could investigate designing lenses whose shape dynamically adapts to the geometry or data of interest, while ensuring proper continuity when transitioning from the focus to the context areas. In this way, these conformal lens shapes would possibly disambiguate the local features in focus, considerably preserve the context, and ultimately reduce cognitive confusion for users when interpreting the lens results.
Effect Imagery The vast majority of the surveyed lenses employ 2D projection images ( out of papers), followed by very few that use 3D () and decal () images. The same pattern holds when considering only immersive environments.
Regarding 3D lens images, as previously mentioned, they may resemble world-in-miniatures whose core affordance lies in serving as both an alternative visualization and interaction space that supplement a large-scale virtual environment—e.g. by selecting an object that is far away from the user’s location. Nonetheless, our survey revealed that volumetric images typically apply the same visual representations as the base visualization. However, there are likely possibilities for further work to investigate novel volumetric lens images that alter the focused visualization in order to depict alternative data facets or visual abstractions—e.g. a lens could apply distortions or perspective cut-offs in its miniature image in order to enable visual access to its partially-occluded focus regions. Beyond that, the world-in-miniature metaphor could be translated as lenses supplementing the base visualization by enabling 3D interactions within the lens imagery itself to adjust lens parameters. Further research could investigate 3D images with distinctive visual abstraction or interaction idioms.
Furthermore, our survey revealed that very little research has investigated combining different categories of lens images ( out of papers). This approach could be particularly beneficial for 3D visualizations, as they often contain a combination of multiple geometric data representations—e.g. points, streamlines, and surfaces, each with associated attributes. In multi-geometry visualizations, users typically need to comprehend (interrelations among) these different spatial data structures in order to gain a holist understanding of the phenomena. In this sense, further research could investigate the design and added benefits of merging different types of lens images, applying each to the data type they are best suited for—e.g. 3D images are suitable for point clouds, while decal-based images best operate on surfaces.
Viewpoint Dependency Viewpoint responsiveness is a heavily underrepresented design strategy, with only papers in our survey featuring it—this holds true in immersive environments. On conventional displays, viewpoint dependency may be overlooked since the limited information space occupies users’ full attention—thus, focus only shifts with explicit interaction. In contrast, on large immersive space, viewpoint changes may signal shifts in user attention and perception of the focus (and context) within the visualization. Further research could consider novel view-sensitive lens designs that adapt their emphasis, effect, function, or even a combination thereof.
Moreover, our survey revealed that immersive lenses are typically placed in close, personal spaces directly in front of users [52, 32, 41]. This is intuitive as the focus is often something that is perceived to be near the viewer. However, in large virtual environments, this may not always hold true—i.e. when users virtually or physically move, the focus of the visualization might still match the center of their attention. To illustrate, in large 3D urban models, continuous perspective changes poses a significant challenge for urban planners when comparing points of interest within the city—e.g. buildings, that are not only distant from each other but may also be occluded by surrounding buildings, depending on the user’s viewpoint [14]. Therefore, future research could investigate design changes in immersive lenses situated outside users’ personal space that still reflect their focus of information.
Furthermore, our survey indicated that the few view-dependent immersive lenses were mainly focused on supplementing individual work, even in collaborative environments—see Figure 11. Therefore, future research could tackle design aspects on perspective-sensitive lenses to support shared work, where multiple users may look at lens images from different points of view. For instance, one could imagine a user communicating focused information to a co-worker situated elsewhere, who has a partially occluded or a relatively narrow viewing angle of the focus region. This could involve rearranging the images to increase its visual angle and perception for the second viewer.
Limitations Lastly, we want to acknowledge that this paper is not a comprehensive literature survey—nor was it intended to be. While our sampled corpus spans work from the past 15 years to reflect a degree of diversity, limitations in sample size and representativeness remain, as the corpus construction mainly relied on manual search and annotation. For that, we refer to an extensive survey conducted by Tominski et al. on lenses at a general level [57].
5 Conclusion
In this work, we tackled the question of what are design components and choices to be considered when constructing spatially-embedded visualization lenses? To this end, we conducted a survey and collected a corpus of 45 papers published in the visualization literature over the past 15 years. From this corpus, we derived a design space comprising seven design dimensions. For each dimension, we distilled information on specific design choices—their usage scenarios, advantages, and disadvantages, supported by a variety of examples. Following this, we described a detailed generative use case encompassing three immersive lenses, thereby conveying the expressive power of our design model.
Finally, guided by our design space, we identified and discussed opportunities for future research on lens design—considering both conventional and emerging immersive environments. For designers, our catalog surfaces a range of design archetypes along with their trade-offs. For researchers, we hope that our work not only stimulates investigations on new designs but also provides a shared vocabulary for how future research understands, describes, and discusses the value of spatially-embedded lens visualizations.

### v2

@@
 Our decision to include immersive lenses was aimed at forming a more comprehensive taxonomy that incorporates design aspects accounting for emerging displays, thereby contributing to the growing field of immersive analytics.
+), Effect Scope (
+), Shape (
+), Dimensionality (
+), Effect Imagery (
+), Effect Encoding (
+), and Viewpoint Dependency (
+) dimensions.
+Some references contain more than one colored cell within a single dimension, given its non-mutually exclusive design choices.
+Papers presenting lens techniques within immersive spaces are marked with an asterisk (
 Methodology To find relevant papers for our review we used a snowball sampling technique starting with the seminal paper by Bier et al.
@@
 Because of the tight interrelation between position, orientation, and scale, we consider them together within a single design dimension. These lens parameters primarily specify the lens selection.
-Most lens techniques supply users full interactivity for setting the lens position, orientation, and scale.
+Most lens techniques supply users full interactivity
+for setting the lens position, orientation, and scale.
 Tominski et. al proposed a circular lens where users interactively set its position and radius, delimiting the lens selection over spatiotemporal data on 2D maps [58].
@@
 By adjusting orientation, the lens can align its geometry with the underlying 3D visualization—e.g. aligning a spatially-tracked mirroring lens with intertwined blood vessels [6].
-Although most surveyed lenses are fully interactive, some rely on (semi-)automated control based on the underlying data [18, 1, 3].
+Although most surveyed lenses are fully interactive, some rely on (semi-)automated
+control based on the underlying data [18, 1, 3].
 Semi-automation may be useful particularly in scenarios where domain knowledge is encoded into the automation logic, guiding users towards relevant data features—e.g. flow patterns.
@@
 The shape of a lens refers to the geometry used for the lens selection, and can be classified as either fixed or dynamic—see Fig. 4.
-Most lenses proposed in the literature are of fixed shape.
+Most lenses proposed in the literature are of fixed
+shape.
 While they typically use simple, primitive forms—e.g.
@@
 However, they may not adequately conform to irregular data structures, resulting in poor coverage of relevant features, inclusion of irrelevant ones, and mental burden as users themselves must disambiguate them.
-Some lens techniques adapt their shape dynamically, either by the user or by the underlying data
+Some lens techniques adapt their shape dynamically, either by the user
+or by the underlying data
 In the former case, lens shapes are typically set by users through sketching, brushing, or lassoing (Fig. 4–right) [48, 11, 24].
@@
 position, orientation, scale, and shape.
-2D lenses are the most commonly used and have polygonal shapes—-e.g. circular or square, which are placed and manipulated in screen space.
+2D lenses
+are the most commonly used and have polygonal shapes—-e.g. circular or square, which are placed and manipulated in screen space.
 For instance, Tim and Elmqvist proposed the use of physical lenses made of thin, transparent paper sheets on top of tabletop surfaces [31].
@@
 Although 2D lenses have been used in 3D visualizations [40, 32, 24, 13], their 2D nature constrains the lens manipulations to the view plane and leads to two primary limitations when handling 3D data: an inability to carry out a 3D selection, and an absence of spatial correlation between the 2D alignment of the lens and the underlying 3D visualization.
-2.5D lenses consist of 2D lenses embedded in 3D spaces, thereby augmenting lens manipulations to 3D [42, 33, 23, 6].
+2.5D lenses
+consist of 2D lenses embedded in 3D spaces, thereby augmenting lens manipulations to 3D [42, 33, 23, 6].
 Although this alleviates the issue of spatial correlation, it increases interaction effort to place and align the lenses according to the underlying data.
@@
 users could arbitrarily slice an MRT scan of a human torso in different poses.
-3D lenses consist of 3D volumes—e.g. a sphere,
+3D lenses
+consist of 3D volumes—e.g. a sphere,
 which enable volumetric spatial selections of the 3D data [28, 61, 9].
@@
 Traditionally, a lens function only applies to visual elements confined inside its shape bounds—-that is, the lens effect scope is the lens
-interior [42, 24, 3].
+interior
+[42, 24, 3].
 Restricting the lens effect to its interior maintains a clear distinction between focus and context.
@@
 Although interior effects offer localized focus+context visualization, it may obscure broader patterns or connections extending beyond the lens boundary—especially in tasks requiring spatial continuitiy, like tracing trajectories.
-A few lens effects impact visual elements located not only inside the lens boundary but also beyond it; therefore, the effect scope is considered to be the lens exterior [41, 50, 19].
+A few lens effects impact visual elements located not only inside the lens boundary but also beyond it; therefore, the effect scope is considered to be the lens exterior
+[41, 50, 19].
 In such cases, the lens boundary is primarily used to define a selection based on a local behavior of interest in the underlying data—e.g.
@@
 They commonly showcase either a duplicate of the visual content enriched with additional elements, or an alternative visual encoding that may not be suitable in the originating view.
-This effect scope is referred to as separate view [46, 10, 50, 27, 58].
+This effect scope is referred to as separate view
+[46, 10, 50, 27, 58].
 Lenses whose effects are displayed across separate views enable side-by-side comparisons; however, careful layout design is needed as to avoid becoming cognitive demanding by requiring users to mentally relate spatially-detached views.
@@
 The effect imagery refers to the visual image generated by the lens function, and can be classified as 2D , 3D, or Decal images—see Fig. 7.
-In 2D lens images the imagery generated as the lens effect consists of projected images, which are rendered either directly onto the view plane [23, 24] or onto a separate render texture to display the lens image as a separate view [42, 46, 6, 5].
+In 2D lens images
+the imagery generated as the lens effect consists of projected images, which are rendered either directly onto the view plane [23, 24] or onto a separate render texture to display the lens image as a separate view [42, 46, 6, 5].
 As an example, Pindat et al.
@@
 In this sense, a view-dependent 2D image may assist restoring the lost spatial perception, as motion parallax effects manifest when projections are continuously rendered in real time [42, 6].
-3D images consist of volumetric visual images, possibly rendered as a separate volume and enriched with additional visual elements [9, 28, 5]—similar to a world-in-miniature [54].
+3D images
+consist of volumetric visual images, possibly rendered as a separate volume and enriched with additional visual elements [9, 28, 5]—similar to a world-in-miniature [54].
 As an example, Johnson et al.
@@
 proposed a data tiling technique to speed up rendering of multiple volumetric images of focus regions on topographic surfaces [9].
-Decal images conform to surface areas of interest.
+Decal images
+conform to surface areas of interest.
 This conforming behavior not only establishes a strong spatial correlation between the lens image and its referent location but also mitigates perceptual issues, such as the lens image itself occluding parts of the base visualization—particularly those of intricate geometry.
@@
 | Lacks spatial correlation | Misaligned, rigid selection | Relies on user decoding | Low spatial perception | ||||
-Prone to clutter |
+| Prone to clutter | |||||||
 | Shape | Viewpoint Dependency | ||||||
@@
 | Data-sensitive fragility | |||||||
-Opaque logic |
+| Opaque logic | |||||||
 | Effect Scope | Position, Orientation & Scale | ||||||
@@
 | Hides broader links | Weaker lens-effect binding | Cross-view burden | Demands spatial dexterity | Misleads if heuristics fail | Fragile on uneven surfaces | ||
-Generalization liability |
+| Generalization liability | |||||||
 2.6 Effect Encoding
 The lens effect enconding can be classified either as thematic or spatially-based, as shown in Fig. 8.
-Thematic encoding draws from information visualization and refers to abstract representations of the lens’ data selection—e.g. linear and radial histograms (Fig. 8–left) [58, 40, 36].
+Thematic encoding
+draws from information visualization and refers to abstract representations of the lens’ data selection—e.g. linear and radial histograms (Fig. 8–left) [58, 40, 36].
 Hurter et al.
@@
 However, physical locomotion can raise visibility and perceptual issues in the focused visualization: visual encodings may become imperceptible with increasing distance (leading to change blindness), distort under acute viewing angles, or even obstruct other foci distributed in the space.
-Although most papers do not consider viewpoint in their designs, view-sensitive lenses afford fluid and intuitive interactions by dynamically adapting to users’ physical navigation or gestures—e.g.
+Although most papers do not consider viewpoint in their designs, view-sensitive
+lenses afford fluid and intuitive interactions by dynamically adapting to users’ physical navigation or gestures—e.g.
 automatically expanding when users move away, and revealing fine-grained information when in close proximity (Fig.